OpenStudy (anonymous):

What is the value of the expression 3–4? A. –81 B. –12 C. –1/81 D. 1/81

OpenStudy (solomonzelman):

\(\Large\color{black}{ \bf 3^{-4}=\frac{1}{3^4}=\frac{\text{?} }{\text{?}} }\)

OpenStudy (anonymous):

1/81

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

You can already exclude all the other options, because no exponent can make a positive number, into a negative number.

OpenStudy (anonymous):

D. 1/81